Drawing three non face cards in a row from a standard deck of cards when the drawn card is not returned to the deck each time what is the probability? How to solve step by step so i can understand i
t. Please and thank you.

Answer by stanbon(75887)

Drawing three non face cards in a row from a standard deck of cards when the drawn card is not returned to the deck each time what is the probability?
-------
There are 3 face cards and 10 non-face cards in each suit
----
# of ways to draw 3 non-face cards: 40C3 = (40*39*38)/(1*2*3) = 9880
----
# of ways to draw 3 cards randomly: 52C3 = 22100
-----
P(draw 3 non-face cards without replacement) = 40C3/52C3 = 0.4471
=============
Cheers,
Stan H.
=============